ppppp:

If sin y° = 9 divided by c and tan y° = 9 divided by d, what is the value of cos y°?

jhonyy9:

sin y = 9/c tan y = 9/d cos y = ? can you rewrite the tan y using sin y and cos y ?

ppppp:

what do you mean

jhonyy9:

what ratio is tan y of sin y and cos y ?

ppppp:

how do i do that

jhonyy9:

do you know the law of sin and the law of cos and the law of tan ?

jhonyy9:

 in case of this triangle sin A = ? using the ratio of sides